OpenStudy (anonymous):

Which provides the most amount of pasta per dollar? A: 3lb pasta for 3.15 B: 4lb pasta for 3.96 C: 5lb pasta for 4.70

Parth (parthkohli):

Divide all to find the unit value. The least amount would be the answer.

OpenStudy (radar):

You could start by making an educated guess. It is obvious that the choice A: is over a dollar per pound, while the choices B: and C: are less than dollar per pound. So you only have to determine which B: or C: provides the lowest unit cost as ParthKohli has suggested. Do you know how to divide to compute the unit cost?

OpenStudy (radar):

To find cost per pound put the cost as numerator and weight (lb) as denominator

OpenStudy (anonymous):

I'd follow what @radar says!

OpenStudy (radar):

Just guessing it looks like choice C: might be the better bargain.

OpenStudy (anonymous):

I guess you are right @radar

OpenStudy (anonymous):

A) 3.15 /3 = 1.05 B) 3.96 /4 = .99 C) 4.70 /5 = .94 So I agree with @radar C is the best answer:)

OpenStudy (radar):

@traile, @jazy, Looks like the guess has been confirmed by jazy. It is the best buy.